Circular No. 7306

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999ev IN NGC 4274 G. M. Hurst, Basingstoke, England, reports the discovery by Tom

Boles, Wellingborough, of a supernova (mag 14.4) on a single unfiltered CCD frame secured for the U.K. Nova/Supernova Patrol on Nov. 7.225 UT with a 0.36-m f/11 reflector. Boles measured the position of SN 1999ev to be α = 12^{h}19^{m}48^{s}.33, δ = +29^{o}37′22′′.2 (equinox 2000.0). M. Armstrong, Rolvenden, provides position end figures 48^{s}.24, 21′′.7 from an unfiltered CCD exposure taken on Nov. 9.211 (mag ∼ 14.5). Hurst notes that nothing is visible at this location on the Vickers Deep Space CCD Atlas: North (limiting mag ∼ 17) or on an image obtained by M. Schwartz (Cottage Grove, OR) on 1998 Dec. 8.

P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum of SN 1999ev was obtained with the Fred L. Whipple Observatory (FLWO) 1.5-m Tillinghast telescope by P. Berlind on Nov. 10.5 UT. The spectrum confirms that the object is a supernova and indicates that it is a type-II event past maximum. A strong Hα P-Cyg feature is present, and the minimum of the absorption gives and expansion velocity of 5300 km/s. Na I and many Fe II features are also present on a continuum that gets brighter toward the red. CCD images obtained by M. Holman with the FLWO 1.2-m telescope at the time of the spectra give a brightness of V = 16.3. A measurement of the position of the supernova yields α = 12^{h}19^{m}48^{s}.31, δ = +29^{o}37′20′′.4 (equinox 2000.0), which is 29′′ west and 28′′ north of the nucleus of NGC 4274.

SATURN VI (TITAN) T. Owen and N. Biver, University of Hawaii; A. Marten, Paris Observa-

tory, Meudon; H. Matthews, Joint Astronomy Centre, Hilo; and R. Meier, Xerox, Zurich, report the detection of the (3-2) line of C^{18}O at 329.33 GHz in the radio spectrum of Titan. The observations were made with the B3 receiver at the James Clerk Maxwell Telescope on Oct. 19.4, 29.4, 31.4, and Nov. 1.4 UT. Concurrent observations of the (3-2) line of CO at 345.80 GHz allowed the use of a model (Hidayat et al. 1998, Icarus 133, 109) to derive a preliminary value of ^{18}O/^{16}O ∼ 2 times the terrestrial value of 2 × 10^{−3}. This is distinctly lower than the enrichment of ^{15}N/^{14}N = 4.5 times terrestrial found in Titan’s HCN (IAUC 6702). The carbon isotopes on Titan exhibit the terrestrial ratio within current uncertainties (Hidayat et al. 1997, Icarus 126, 170).

1999 November 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green